Sol Summary – November 10th

Sol: 11

Summary Title: Winter is coming to Mars

Author’s Name: Cristina Vazquez Reynel, Crew Exec Officer and HSO

Mission Status: Nominal

Sol Activity Summary: Today the crew woke up at 8am full of energy and had a nice breakfast together.

The crew had a busy day ahead with Laurene, Paula and Marta having braided sessions with their friends and family and two EVAs planned.

The day started by preparing for the EVA experiments and cleaning in the RAM and GreenHab. Just before lunch, Laurene and Cristina went out for the first EVA of the day, they tested the robot close to the Hab and went to explore the amazing views of Skyline Rim. Unfortunately, the test failed and no useful data was collected this time. Following their return to the Hab, Paula and Marta got ready and went out for the second EVA, they successfully repeated the test with the robot and spent some time exploring and taking nice pictures with the beautiful sunset.

After the cold EVAs, with temperatures below 0°C, the crew gathered for a hot chocolate and started preparing the reports of the day.

Look ahead plan: Tomorrow the crew is planning a full day preparing for the end of the simulation and a special trip to the Overlook.

Anomalies at work: N/A

Weather: Sunny, 3°C

Crew Physical Status: Good

EVA: 2

Reports to be filed: Sol Summary Report, Commander Report, GreenHab Report, Operations Report, EVA Request

Support Requested: None, thank you Mission Support!

Operations Report – November 10th

SOL: 11
Name of person filing report: Laurène Delsupexhe
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Spirit rover used: No
Hours: (before EVA) N/A
Beginning charge: (Before EVA) N/A
Ending charge: (On return from EVA, before recharging) N/A
Currently charging: N/A
Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A
General notes and comments: N/A
Summary of Hab operations:
Water (static tank): 258.0 gallons
Water (loft tank): N/A
Water Meter: N/A
Static to Loft Pump used – no
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes
Summary of internet: Nothing to report
Summary of suits and radios: Nothing to report
Summary of GreenHab operations:
WATER USE: 7.1 gallons
Heater: on
Supplemental light: off
Harvest: yes
Summary of Science Dome operations:
Dual split: Heat or AC, Off
Summary of RAM operations: N/A
Summary of any observatory issues: N/A
Summary of health and safety issues: N/A
Questions, concerns and requests to Mission Support: N/A

GreenHab Report – November 10th

GreenHab Officer: Paula Peixoto Moledo
Environmental control: 17% Humidity
Average temperatures: 27,5ºC/ 28ºC
Hours of supplemental light: N/A
Daily water usage for crops: 27L ≈ 7.1 Gallons
Daily water usage for research and/or other purposes: None
Water in Blue Tank 35.8 gallons
Time(s) of watering for crops: 9:30 am – 7:00 pm
Changes to crops:
Narrative: I watered the plants 2 times today as every day since the beginning of the mission. We cleaned and tidied up a little bit of the GreenHab today ! As good as new !
Support/supplies needed: -.

EVA Report – November 10th

EVA # 13
Author: Laurène Delsupexhe
Purpose of EVA: Experiment with robot Lazarus + Exploring Sagan Street
Start time: 11:50am
End time: 1:40 pm
Narrative: The crew left the main Hab at 11:50am with the robot and tested 3 times the same straight path of 10m close to the observatory. After that, they left the equipment in the airlock for the team in the Hab to download the data and recharge the battery. Meanwhile, Laurène and Cristina headed towards the North Ridge in order to reach Sagan Street. The Mission Support Crew contacted them to let them know the testing of the robot had not properly worked. As there was still an hour and a half in the EVA and the crew was already halfway to Sagan Street, they decided to continue the exploration part of the EVA. They reached their destination, took some samples to test later on and headed back to the station by the same way they arrived. They ended their EVA as soon as they came back to the station and decided that Paula and Marta will redo the test during their EVA.
Destination: Sagan Street
Coordinates (use UTM NAD27 CONUS): 4251500-516500
Participants: Cristina Vazquez and Laurène Delsupexhe
Road(s) and routes per MDRS Map: Sagan Street, 1103
Mode of travel: Walking

EVA #: 14
Author: Marta Ferran
Purpose of EVA: Experiment with robot Lazarus + Exploring
Start time: 2:40pm
End time: 5:30pm
Narrative: Paula and Marta left the main Hab at 2:30 PM with Lazarus, the visual odometry robot developed by Dronomy. The crew tested 4 times the same straight path of 12 m close to the MDRS. Once finished, the crew came back to the Hab and left the robot equipment in the airlock. After that, Paula and Marta continued with their exploration around the MDRS. The sky was gorgeous and so, the crew took amazing pictures of the red planet. After almost 3h of EVA, the crew headed back to the Hab and started working on Lazarus’ files.
Destination: Around MDRS
Coordinates (use UTM NAD27 CONUS): 4251000-518000
Participants: Marta Ferran and Paula Peixoto
Road(s) and routes per MDRS Map: N/A
Mode of travel: Walking

Commander Report – November 10th>

Sol: 11
Summary Title: Freezing Mars
Author’s name: Marta Ferran-Marqués, Commander and Crew Scientist
Mission Status: Nominal
Commander Report: The WoMars crew woke up this morning at 8 am. After the commander’s debrief, Laurene and Cristina went to the RAM to clean everything up. Meanwhile, Paula and Marta went to the GreenHab to water up the plants and clean the space. Later in the morning, Laurene and Paula had their last time-delayed communication sessions with their loved ones.

Cristina and Laurene went for an EVA around the surroundings of the MDRS to continue with Lazarus’ experiments – the visual odometry robot developed by Dronomy. After lunch, Paula and Marta also went for an EVA around the surroundings of the MDRS and completed the experiments with Lazarus for the Nerio I mission. The sky was gorgeous today, showing blue, pink and purple colors.

Once back in the Hab, Marta had her last time-delayed communication session with her brother. After writing all the reports, the crew will cook dinner and have a relaxing night!

Copyright © The Mars Society. All rights reserved. | Main Site